# Projective Modules and Cohomology for Integral Basic Algebras

### v1

@@ Abstract.
Algebras defined over fields of characteristic zero and positive characteristic usually do not behave the same way. However, for certain algebras, for example the group algebras, they behave the same way as the characteristic zero case at good enough prime. In this paper, we initiate the study of this topic by imposing increasingly strong hypotheses on basic algebras. When the algebras satisfy the right hypotheses, we have equalities of the dimensions of their cohomology groups between simples modules and equalities of graded Cartan numbers. The examples include the Solomon descent algebras of finite Coxeter groups at large enough primes, nil-Coxeter algebra, and certain finite semigroup algebras at an arbitrary prime.
@@ 1. Introduction
In the representation theory of finite groups, characteristic zero and good enough prime characteristic behave the same way. Namely in both cases, the group algebra is semisimple. If the fields are algebraically closed then the simple modules correspond and have the same dimensions. Here, the characteristic is good enough if it does not divide the group order.
@@
There are various questions that arise here, and correspondingly there are various notions of good enough. For example, one might at least want the simple modules to correspond, and have the same dimensions. And then, how similar are the cohomology algebras? Of course, they cannot be isomorphic, as they have different characteristics. But at least we can compare their dimensions in each degree. To refine this, we could examine the product structure and secondary operations, define various constants in terms of these, and compare them. We can also compare the Loewy layers of their projective indecomposable modules, and other more subtle invariants of this sort.
@@
The examples of basic algebras we examine in detailed are the Solomon descent algebra of a finite Coxeter group [1, 2, 3, 6, 9, 10, 11, 12, 15, 18, 23, 27, 29, 31], the nil-Coxeter algebra [7, 16, 21], and certain classes of finite semigroup algebras [5, 8, 24, 25, 28, 30]. The descent algebra was originally constructed by Solomon as a non-commutative analogue of the Mackey’s multiplication formula for permutation characters of the parabolic subgroups of the Coxeter group. But it can also be obtained as the opposite algebra of the fixed-point space of a certain face algebra of hyperplane arrangement in real space associated with the Coxeter group. The face semigroup algebra belongs to a broader class known as unital left regular band algebras, which, in turn, are part of the even larger class known as the $\mathcal{R}$-trivial monoid algebras. Notably, this largest class also includes the 00-Hecke algebras.
We show that the strongest of our hypotheses holds for the descent algebra for sufficiently large primes and nil-Coxeter algebra for an arbitrary prime. In the case of the semigroup algebras that we investigate, we show that the weaker hypotheses hold under suitable conditions, but we leave open the question of when the stronger one does. While our results do not give an explicit description of the cohomology groups nor the multiplicities of simple modules in Loewy layers of projective indecomposables for these classes of algebras, they offer a general explanation why these notions behave the same way in characteristic zero and at good enough prime.
In preparation for this task, in the next section, we give a general discussion of various hypotheses (Hypotheses AE) on a finitely generated algebra over a local principal ideal domain, and what they imply. The main results are summarised in Theorem 2.3. Sections 3, 4, 5 and 6 are devoted to the proof of the theorem. Then we apply the results to the cases of the descent algebra of a finite Coxeter group, nil-Coxeter algebra, and certain semigroup algebras and obtain Theorems 9.15, 11.2, 12.2 and 13.2. We begin the general discussion on the descent algebra in Section 7 and extend the result of the second author regarding the idempotents of the descent algebra to all types and all primes in Section 8. In Section 9, we show that the descent algebra satisfies our strongest of our hypotheses at large enough prime. The analysis is based on a certain characteristic free basis of the descent algebra and we further investigate that in Section 10. Another algebra satisfying our strongest hypothesis is the nil-Coxeter algebra and that is discussed in Section 11. In Sections 12 and 13, we show that certain classes of semigroup algebras satisfy our weaker hypotheses.
@@ 3. Quivers and relations over ${\mathcal{O}}$
Here are a couple of examples that do not satisfy D. The first is $A={\mathcal{O}}[x]/(x^{2}-x^{3})$. In this case, $Q$ has just one vertex, it has one loop corresponding to $x$, but $I=(x^{2}-x^{3})$ contains no power of $J_{Q}=(x)\leqslant{\mathcal{O}}Q$. So $J$ is not a nilpotent ideal in $A$. The second example is $A={\mathcal{O}}[x]/(\pi x^{2},x^{3})$. Then $x^{2}\neq 0$ but $\pi x^{2}=0$ in $A$, so $A$ is not ${\mathcal{O}}$-free. In this example we have $\dim_{\mathbb{K}}{\hat{A}}=2$, $\dim_{\mathbb{F}}{\bar{A}}=3$, so part (v) of Theorem 3.1 does not hold.
@@ 4. ${\mathcal{O}}$-freeness of $\operatorname{\sf Ext}$
Since $\pi A\subseteq\operatorname{\sf Rad}(A)$, $A/\operatorname{\sf Rad}(A)$ is a finite dimensional semisimple ${\mathbb{F}}$-algebra and idempotents lift to $A$, it follows that $A$ is semiperfect. See Lam [22, Chapter 8], especially Definition 23.1 and Propositions 23.5, 24.10 and 24.12. ∎
@@ 5. Comparing $\operatorname{\sf Ext}$ over ${\mathcal{O}}$, ${\mathbb{F}}$ and $\mathbb{K}$
We have $\operatorname{\sf Tor}_{s}^{A}(A,M)=0$ for $s\geqslant 1$, so $\operatorname{\sf Tor}^{A}_{s}({\bar{A}},M)=0$ for $s\geqslant 2$. Since $M$ is ${\mathcal{O}}$-torsion free and multiplication with $\pi$ is injective, it also follows that $\operatorname{\sf Tor}^{A}_{1}({\bar{A}},M)=0$. By Cartan and Eilenberg [14, VI, Proposition 4.1.3], we get an isomorphism
@@ 6. ${\mathcal{O}}$-freeness of $J^{n}/J^{n+1}$
and $A={\mathcal{O}}Q/I$ with $I=(edc-\pi ba)$. This satisfies D, but not E, because $ba$ is in $J^{2}$ but not in $J^{3}$ while $\pi ba$ is in $J^{3}$. So as an ${\mathcal{O}}$-module, $J^{2}/J^{3}\cong{\mathbb{F}}\oplus{\mathcal{O}}\oplus{\mathcal{O}}$, spanned by $ba$, $dc$ and $ed$. Thus $A/J^{3}\cong{\mathbb{F}}\oplus 7.{\mathcal{O}}$ is not ${\mathcal{O}}$-free.
@@ 7. The descent algebras
In this section, we review some elementary properties for (finite) Coxeter groups and their descent algebras for the discussions in Sections 8, 9 and 10. For the general knowledge about Coxeter groups, we refer the reader to [19, 20].
@@
Theorem 7.3 ([31, Theorem 3], [2, Theorem 3]).
@@
For example, if $n=1$ and $J=_{W}K$, we have $x_{(J,K)}=x_{J}-x_{K}\in\operatorname{\sf Rad}(\mathscr{D}_{k})$. By Theorem 7.3, if $q\nmid|W|$, we have
@@
Lemma 7.5 ([19, Proposition 2.4.4] and [2, Lemma 1]).
@@ 8. Idempotents for the descent algebras
The proof of the following lemma is similar to that of [23, Lemma 3.2] and we shall leave it to the readers.
@@
Since we are dealing with the characteristic zero case ($q=0$) as well, we need to replace the idempotent-lifting procedure in [23, Definition 3.5] with the $(3a^{2}-2a^{3})$-construction and provide proofs for both Lemma 8.6 and Theorem 8.7. For this, we need to introduce some extra notation as follows.
@@
We refer the readers to the proof of [4, Theorem 1.7.3]. The fact that $a_{i}$ commutes with $a_{1}$ can be proved easily using induction on $i$. Notice that $(a_{1}a_{\infty})+N$ is again an idempotent. Suppose that $b_{1}=a_{1}a_{\infty}$. We check that $b_{i}=a_{i}a_{\infty}$ for all $i\geqslant 1$ by induction. Assume that $a_{i}$ commutes with $a_{1}$ and since $a_{\infty}$ is an idempotent, we have
@@
The following lemma is similar to [23, Lemma 3.6].
@@
Lemma 8.6 and [23, Proposition 3.3(ii)] give us the first sentence of our theorem. We now prove inductively on $i$ that $e_{J}$ has the desired form. When $i=1$, since $Y^{\circ}_{\leqslant 1}=Y^{\circ}_{J_{1}}$, we have
@@ 9.1. Verifying B
The proof of the following proposition requires the construction presented in Section 8 which we shall not go into the details.
@@ 9.2. Verifying E
Part (ii) follows from Lemma 7.4. We now prove part (i). For each $(\underline{J},\underline{K})\in{\mathsf{J}}(n)$ and $L\subseteq S$, since the structure constants of the descent algebra belong in ${\mathbb{Z}}$, we have $x_{\underline{J},\underline{K}}\in\mathscr{D}_{\mathbb{Z}}$. Also, since $x_{\underline{J},\underline{K}}\in\operatorname{\sf Rad}(\mathscr{D}_{\mathbb{Q}})=\mathrm{span}_{\mathbb{Q}}\mathcal{B}$, the sum of all the coefficients is 00 and hence the entries in $[{\mathsf{J}}(n)]$ belong in ${\mathbb{Z}}$. ∎
@@
Let $(W,S)$ be a Coxeter system and $s=\dim_{\mathbb{K}}\operatorname{\sf Rad}^{n}({\hat{\mathscr{D}}})\neq 0$ (or equivalently, $[{\mathsf{J}}(n)]$ is not a zero matrix). Let $d_{W,n}$ be the greatest common divisor of the determinants of all possible non-singular $(s\times s)$-submatrices of $[{\mathsf{J}}(n)]$. Furthermore, let $n_{W,n}=\operatorname{\sf lcm}(d_{W,n},|W|)$ and $n_{W}$ be the least common multiple of the $n_{W,n}$’s such that $d_{W,n}\neq 0$, i.e.,
@@
Consider the Coxeter group $(W,S)$ of the dihedral type $\mathbb{I}_{n}$ where $S=\{s_{1},s_{2}\}$ so that $W$ is the dihedral group of order $2n$. It is well-known that $\{s_{1}\}=_{W}\{s_{2}\}$ if and only if $n$ is odd. Therefore, $\Omega=\emptyset$ unless $n$ is odd. In the case when $n$ is odd, $B=\{x_{\{s_{1}\}}-x_{\{s_{2}\}}\}$ and $[{\mathsf{J}}(n)]$ is zero for $r\geqslant 2$. Therefore, for each $n$, we have $n_{W}=2n$.
@@
The matrix $[{\mathsf{J}}(n)]$ is zero for $r\geqslant 3$. For $r=2$, $\operatorname{\sf Rad}^{2}({\hat{\mathscr{D}}})$ is one-dimensional. Since
@@
By Magma [13] computation, for type $\mathbb{E}_{6}$ or $\mathbb{E}_{7}$ and for each $n\in\{2,3,4\}$, we found a submatrix of $[{\mathsf{J}}(n)]$ with the prime factors of whose determinant are only $2,3,5$. For example, in the case of $\mathbb{E}_{7}$, one of the determinants is a 53-digit number but with only prime factors $2$, $3$ and $5$.
@@ 10. Independence of choice of basis for $n_{W}$
11. Nil-Coxeter Algebras satisfy E
In this section, we show that the nil-Coxeter algebra satisfies our strongest hypothesis. The algebra originates in [7, Theorem 3.4]. Subsequently, for example, it has been used in the study of the Schubert polynomials in [16] and studied for its categorification aspect in [21].
Fix a finite Coxeter system $(W,S)$. For any integral domain $R$, we define the nil-Coxeter algebra $\mathscr{N}_{R}(W)$ as the $R$-algebra generated by $\{x_{s}:s\in S\}$ subject to the relations $x_{s}^{2}=0$ and
@@ 11. Nil-Coxeter Algebras satisfy E
Conjecture 11.3.
Let $T$ be the unique simple $\mathscr{N}_{k}(W)$-module and $n$ be the rank of the Coxeter system $(W,S)$. The Hilbert–Poincaré series for the Ext algebra $\operatorname{\sf Ext}^{*}_{\mathscr{N}_{k}(W)}(T,T)$ is
$\sum_{i=0}^{\infty}t^{i}\left(\dim_{k}\operatorname{\sf Ext}^{i}_{\mathscr{N}_{k}(W)}(T,T)\right)=\frac{1}{(1-t)^{n-1}}.$

### v2

@@ Abstract.
Algebras defined over fields of characteristic zero and positive characteristic usually do not behave the same way. However, for certain algebras, for example the group algebras, they behave the same way as the characteristic zero case at “good enough” prime. In this paper, we initiate the study of this topic by imposing increasingly strong hypotheses on basic algebras. When the algebras satisfy the right hypotheses, we have equalities of the dimensions of their cohomology groups between simple modules and equalities of graded Cartan numbers. The examples include the Solomon descent algebras of finite Coxeter groups at large enough primes, nilCoxeter algebra, and certain finite semigroup algebras at an arbitrary prime.
@@ 1. Introduction
In the representation theory of finite groups, characteristic zero and “good enough” prime characteristic behave the same way. Namely in both cases, the group algebra is semisimple. If the fields are algebraically closed then the simple modules correspond and have the same dimensions. Here, the characteristic is good enough if it does not divide the group order.
@@
There are various questions that arise here, and correspondingly there are various notions of “good enough”. For example, one might at least want the simple modules to correspond, and have the same dimensions. And then, how similar are the cohomology algebras? Of course, they cannot be isomorphic, as they have different characteristics. But at least we can compare their dimensions in each degree. To refine this, we could examine the product structure and secondary operations, define various constants in terms of these, and compare them. We can also compare the Loewy layers of their projective indecomposable modules, and other more subtle invariants of this sort.
@@
The examples of basic algebras we examine in detailed are the Solomon descent algebra of a finite Coxeter group [1, 2, 3, 7, 10, 11, 12, 13, 16, 19, 24, 28, 29, 31, 33], the nilCoxeter algebra [8, 17, 22], and certain classes of finite semigroup algebras [6, 9, 25, 26, 30, 32]. The descent algebra was originally constructed by Solomon as a non-commutative analogue of the Mackey’s multiplication formula for permutation characters of the parabolic subgroups of the Coxeter group. But it can also be obtained as the opposite algebra of the fixed-point space of a certain face algebra of hyperplane arrangement in real space associated with the Coxeter group. The face semigroup algebra belongs to a broader class known as unital left regular band algebras, which, in turn, are part of the even larger class known as the $\mathcal{R}$-trivial monoid algebras. Notably, this largest class also includes the 00-Hecke algebras.
We show that the strongest of our hypotheses holds for the descent algebra for sufficiently large primes and the nilCoxeter algebra for an arbitrary prime. In the case of the semigroup algebras that we investigate, we show that the weaker hypotheses hold under suitable conditions, but we leave open the question of when the stronger one does. While our results do not give an explicit description of the cohomology groups nor the multiplicities of simple modules in Loewy layers of projective indecomposables for these classes of algebras, they offer a general explanation why these notions behave the same way in characteristic zero and at “good enough” prime.
In preparation for this task, in the next section, we give a general discussion of various hypotheses (Hypotheses AE) on a finitely generated algebra over a local principal ideal domain, and what they imply. The main results are summarised in Theorem 2.3. Sections 3, 4, 5 and 6 are devoted to the proof of the theorem. Then we apply the results to the cases of the descent algebra of a finite Coxeter group, the nilCoxeter algebra, and certain semigroup algebras and obtain Theorems 9.15, 11.2, 12.2 and 13.2. We begin the general discussion on the descent algebra in Section 7 and extend the result of the second author regarding the idempotents of the descent algebra to all types and all primes in Section 8. In Section 9, we show that the descent algebra satisfies our strongest of our hypotheses at large enough prime. The analysis is based on a certain characteristic free basis of the descent algebra and we further investigate that in Section 10. Another algebra satisfying our strongest hypothesis is the nilCoxeter algebra and that is discussed in Section 11. In Sections 12 and 13, we show that certain classes of semigroup algebras satisfy our weaker hypotheses.
@@ 3. Quivers and relations over ${\mathcal{O}}$
Here are a couple of examples that do not satisfy D. The first is $A={\mathcal{O}}[x]/(x^{2}-x^{3})$. In this case, $Q$ has just one vertex, it has one loop corresponding to $x$, but $I=(x^{2}-x^{3})$ contains no power of $J_{Q}=(x)\leqslant{\mathcal{O}}Q$. So $J_{A}$ is not a nilpotent ideal in $A$. The second example is $A={\mathcal{O}}[x]/(\pi x^{2},x^{3})$. Then $x^{2}\neq 0$ but $\pi x^{2}=0$ in $A$, so $A$ is not ${\mathcal{O}}$-free. In this example we have $\dim_{\mathbb{K}}{\hat{A}}=2$, $\dim_{\mathbb{F}}{\bar{A}}=3$, so part (v) of Theorem 3.1 does not hold.
@@ 4. ${\mathcal{O}}$-freeness of $\operatorname{\sf Ext}$
Since $\pi A\subseteq\operatorname{\sf Rad}(A)$, $A/\operatorname{\sf Rad}(A)$ is a finite dimensional semisimple ${\mathbb{F}}$-algebra and idempotents lift to $A$, it follows that $A$ is semiperfect. See Lam [23, Chapter 8], especially Definition 23.1 and Propositions 23.5, 24.10 and 24.12. ∎
@@ 5. Comparing $\operatorname{\sf Ext}$ over ${\mathcal{O}}$, ${\mathbb{F}}$ and $\mathbb{K}$
We have $\operatorname{\sf Tor}_{s}^{A}(A,M)=0$ for $s\geqslant 1$, so $\operatorname{\sf Tor}^{A}_{s}({\bar{A}},M)=0$ for $s\geqslant 2$. Since $M$ is ${\mathcal{O}}$-torsion free and multiplication with $\pi$ is injective, it also follows that $\operatorname{\sf Tor}^{A}_{1}({\bar{A}},M)=0$. By Cartan and Eilenberg [15, VI, Proposition 4.1.3], we get an isomorphism
@@ 6. ${\mathcal{O}}$-freeness of $J^{n}/J^{n+1}$
and $A={\mathcal{O}}Q/I$ with $I=(edc-\pi ba)$. Let $J=J_{A}$. The algebra $A$ satisfies D, but not E, because $ba$ is in $J^{2}$ but not in $J^{3}$ while $\pi ba$ is in $J^{3}$. So as an ${\mathcal{O}}$-module, $J^{2}/J^{3}\cong{\mathbb{F}}\oplus{\mathcal{O}}\oplus{\mathcal{O}}$, spanned by $ba$, $dc$ and $ed$. Thus $A/J^{3}\cong{\mathbb{F}}\oplus{\mathcal{O}}^{\oplus 12}$ is not ${\mathcal{O}}$-free.
@@ 7. The descent algebras
In this section, we review some elementary properties for (finite) Coxeter groups and their descent algebras for the discussions in Sections 8, 9 and 10. For the general knowledge about Coxeter groups, we refer the reader to [20, 21]. We will investigate the algebra over a general field $k$ of characteristic $q$ (see 2.1).
@@
Theorem 7.3 ([33, Theorem 3], [2, Theorem 3]).
@@
For example, if $n=1$ and $J=_{W}K$, we have $x_{(J),(K)}=x_{J}-x_{K}\in\operatorname{\sf Rad}(\mathscr{D}_{k})$. By Theorem 7.3, if $q\nmid|W|$, we have
@@
Lemma 7.5 ([20, Proposition 2.4.4] and [2, Lemma 1]).
@@ 8. Idempotents for the descent algebras
The proof of the following lemma is similar to that of [24, Lemma 3.2] and we shall leave it to the readers.
@@
Since we are dealing with the characteristic zero case ($q=0$) as well, we need to replace the idempotent-lifting procedure in [24, Definition 3.5] with the $(3a^{2}-2a^{3})$-construction (see the proof of [5, Theorem 1.7.3]) and provide proofs for both Lemma 8.6 and Theorem 8.7. For this, we need to introduce some extra notation as follows.
@@
The first assertion is well-known. The fact that $a_{i}$ commutes with $a_{1}$ can be proved easily using induction on $i$. Notice that $(a_{1}a_{\infty})+N$ is again an idempotent. To show that $(a_{1}a_{\infty})_{\infty}=a_{\infty}$, let $b_{1}=a_{1}a_{\infty}$ and, inductively, for all $i\geqslant 1$, let $b_{i+1}=3b_{i}^{2}-2b_{i}^{3}$. We check that $b_{i}=a_{i}a_{\infty}$ for all $i\geqslant 1$ using induction. Since $a_{\infty}$ is an idempotent, we have
@@
The following lemma is similar to [24, Lemma 3.6].
@@
Lemma 8.6 and [24, Proposition 3.3(ii)] give us the first sentence of our theorem. We now prove inductively on $i$ that $e_{J}$ has the desired form. When $i=1$, since $Y^{\circ}_{\leqslant 1}=Y^{\circ}_{J_{1}}$, we have
@@ 9.1. Verifying B
The proof of the following lemma requires the construction presented in Section 8 which we shall not go into the details.
@@ 9.2. Verifying E
Part (ii) follows from Lemma 7.4. We now prove part (i). For each $(\underline{J},\underline{K})\in{\mathsf{J}}(n)$ and $L\subseteq S$, since the structure constants of the descent algebra belong in ${\mathbb{Z}}$, we have $x_{\underline{J},\underline{K}}\in\mathscr{D}_{\mathbb{Z}}$. Also, since $x_{\underline{J},\underline{K}}\in\operatorname{\sf Rad}(\mathscr{D}_{\mathbb{Q}})=\mathrm{span}_{\mathbb{Q}}\mathcal{B}$, the sum of all the coefficients in terms of the basis $\{x_{J}:J\subseteq S\}$ is 00 and hence the entries in $[{\mathsf{J}}(n)]$ belong in ${\mathbb{Z}}$. ∎
@@
Let $(W,S)$ be a Coxeter system and $s=\dim_{\mathbb{K}}\operatorname{\sf Rad}^{n}({\hat{\mathscr{D}}})\neq 0$ (or equivalently, $[{\mathsf{J}}(n)]$ is not a zero matrix). Let $d_{W,n}$ be the greatest common divisor of the determinants of all possible $(s\times s)$-submatrices of $[{\mathsf{J}}(n)]$. Furthermore, let $n_{W,n}=\operatorname{\sf lcm}(d_{W,n},|W|)$ and $n_{W}$ be the least common multiple of the $n_{W,n}$’s such that $d_{W,n}\neq 0$, i.e.,
@@
Consider the Coxeter group $(W,S)$ of the dihedral type $\mathbb{I}_{n}$ where $S=\{s_{1},s_{2}\}$ so that $W$ is the dihedral group of order $2n$. It is well-known that $\{s_{1}\}=_{W}\{s_{2}\}$ if and only if $n$ is odd. Therefore, $\Omega=\emptyset$ unless $n$ is odd. In the case when $n$ is odd, $B=\{x_{\{s_{1}\}}-x_{\{s_{2}\}}\}$ and $[{\mathsf{J}}(r)]$ is zero for $r\geqslant 2$. Therefore, for each $n$, we have $n_{W}=2n$.
@@
The matrix $[{\mathsf{J}}(r)]$ is zero for $r\geqslant 3$. For $r=2$, $\operatorname{\sf Rad}^{2}({\hat{\mathscr{D}}})$ is one-dimensional. Since
@@
By Magma [14] computation, for type $\mathbb{E}_{6}$ or $\mathbb{E}_{7}$ and for each $n\in\{2,3,4\}$, we found a submatrix of $[{\mathsf{J}}(n)]$ with the prime factors of whose determinant are only $2,3,5$. For example, in the case of $\mathbb{E}_{7}$, one of the determinants is the following 53-digit number:
[…]
@@ 10. Independence of choice of basis for $n_{W}$
11. nilCoxeter Algebras satisfy E
In this section, we show that the nilCoxeter algebra satisfies our strongest hypothesis. The algebra originates in [8, Theorem 3.4]. Subsequently, for example, it has been used in the study of the Schubert polynomials in [17] and studied for its categorification aspect in [22].
Fix a finite Coxeter system $(W,S)$. For any integral domain $R$, we define the nilCoxeter algebra $\mathscr{N}_{R}(W)$ as the $R$-algebra generated by $\{x_{s}:s\in S\}$ subject to the relations $x_{s}^{2}=0$ and
@@
Let $T$ be the unique simple $\mathscr{N}_{k}(W)$-module. Theorems 11.2 and 2.3 imply that both $\dim_{k}\operatorname{\sf Ext}^{i}_{\mathscr{N}_{k}(W)}(T,T)$ and hence the Hilbert–Poincaré series for the Ext algebra $\operatorname{\sf Ext}^{*}_{\mathscr{N}_{k}(W)}(T,T)$ are independent of the field $k$. In the recent article [4], the first author studied the type $\mathbb{A}$ case. He computed the Hilbert–Poincaré series and presented a set of generators and relations for the Ext algebra. We refer the readers to the article for the full statement.
Theorem 11.3 ([[4, Theorem 1.1]).
Suppose that $n\geqslant 2$. The Ext algebra $\operatorname{\sf Ext}^{*}_{\mathscr{N}_{\mathbb{Z}}(\mathfrak{S}_{n})}({\mathbb{Z}},{\mathbb{Z}})$ has the Hilbert–Poincaré series
$\sum_{i=0}^{\infty}t^{i}\left(\mathsf{rank}_{\mathbb{Z}}\operatorname{\sf Ext}^{i}_{\mathscr{N}_{\mathbb{Z}}(\mathfrak{S}_{n})}({\mathbb{Z}},{\mathbb{Z}})\right)=\frac{1}{(1-t)^{n-1}}.$